OpenStudy (anonymous):

Solve the system of equations by substitution. 7x-9y=-48 -7x-7y=0

OpenStudy (anonymous):

7x-9y=-48 -7x-7y=0 how do I say this... Do you want me just to write it out how to solve it?

OpenStudy (anonymous):

yes